# Watson-Crick Conjugates of Words and Languages

### v3

@@ 5. Iterated $\theta$-conjugate of Words And Languages
Explanation of Step 2 through an Example: After Step 1, we have $w=C_{a_{1}}C_{b_{1}}C_{a_{1}}C_{\theta(a_{2})}$. Then, $S=\{C_{a_{1}}C_{b_{1}}C_{a_{1}}C_{\theta(a_{2})},C_{a_{1}}C_{b_{1}}C_{\theta(a_{2})}C_{a_{1}},C_{a_{1}}C_{a_{1}}C_{b_{1}}C_{\theta(a_{2})},C_{a_{1}}C_{a_{1}}C_{\theta(a_{2})}C_{b_{1}},C_{a_{1}}C_{\theta(a_{2})}C_{a_{1}}C_{b_{1}},C_{a_{1}}C_{\theta(a_{2})}C_{b_{1}}C_{a_{1}},\\ C_{b_{1}}C_{a_{1}}C_{a_{1}}C_{\theta(a_{2})},C_{b_{1}}C_{a_{1}}C_{\theta(a_{2})}C_{a_{1}},C_{b_{1}}C_{\theta(a_{2})}C_{a_{1}}C_{a_{1}},C_{\theta(a_{2})}C_{a_{1}}C_{a_{1}}C_{b_{1}},C_{\theta(a_{2})}C_{a_{1}}C_{b_{1}}C_{a_{1}},C_{\theta(a_{2})}C_{b_{1}}C_{a_{1}}C_{a_{1}}\}$. Using multinomial coefficients, $|S|={4\choose 2~{}1~{}1}=12$.
@@
We now show that $R=C_{\theta}^{*}(w)$. Consider $u\in R$. Then by the above construction, $|u|_{a,\theta(a)}=|w|_{a,\theta(a)}$ for all $a\in\Sigma$. Thus, $u\in C_{\theta}^{*}(w)$. Therefore, $R\subseteq C_{\theta}^{*}(w)$. Consider $v\in C_{\theta}^{*}(w)$. Then, $|v|_{a,\theta(a)}=|w|_{a,\theta(a)}$ for all $a\in\Sigma$. Now, we replace each $a_{i}$ and $\theta(a_{i})$ in $v$ and $w$ by $C_{a_{i}}$ where $1\leq i\leq l$. Also, we replace each $b_{j}$ in $v$ and $w$ by $C_{b_{j}}$ where $1\leq j\leq r$. After replacements, consider $v$ and $w$ become $v^{\prime}$ and $w^{\prime}$, respectively, over $C_{1}\cup D$. Since $|v|_{a,\theta(a)}=|w|_{a,\theta(a)}$ for all $a\in\Sigma$, $v^{\prime}$ is a permutation of $w^{\prime}$, i.e., $v^{\prime}\in S$. Then by step 3, $v\in R$. Thus, $C_{\theta}^{*}(w)\subseteq R$. Therefore, $R=C_{\theta}^{*}(w)$.
@@
  1. (1)
[…]
  2. (2)
[…]
  3. (3)
[…]
@@
It can be observed from Remark 5.11 that if $L$ is finite, then ${C_{\theta}}^{i}(L)$ is finite for $i\geq 0$. Also, if ${C_{\theta}}^{i}(L)$ is finite for some $i\geq 0$, then as $L\subseteq{C_{\theta}}^{i}(L)$, $L$ is finite. This implies that $L$ is finite if and only if ${C_{\theta}}^{i}(L)$ is finite for $i\geq 0$.
@@
Consider a regular language $L=\{(bce)^{*}\}$ over $\{b,c,d,e,f\}$ where $\theta(b)=b$, $\theta(c)=d$ and $\theta(e)=f$. Since ${C_{\theta}}^{*}(w)=\{u\;|\;|u|_{a,\theta(a)}=|w|_{a,\theta(a)}~{}\forall a\in\Sigma\}$, ${C_{\theta}}^{*}(L)=\bigcup\limits_{w\in L}\{u\;|\;|u|_{a,\theta(a)}=|w|_{a,\theta(a)}~{}\forall a\in\Sigma\}.$ Then, ${C_{\theta}}^{*}(L)\cap\{b^{m_{1}}c^{m_{2}}e^{m_{3}}\;|\;m_{i}\geq 0\}=\{b^{n}c^{n}e^{n}\;|\;n\geq 0\}$, which is not context-free (using Theorem 4.4). Note that $\{b^{m_{1}}c^{m_{2}}e^{m_{3}}\;|\;m_{i}\geq 0\}$ is a regular language. Since the regular languages and the context-free languages are closed under intersection with regular languages ([35]), the claim holds. ∎
6. $\theta$-conjugate-free Languages
@@
6.1. Decidability of $\theta$-conjugate-freeness
@@
Define an empty array $A$ of size $n$
@@
     for $l=n$ to $1$ do $z\leftarrow z~{}||~{}\theta(u[l])$triangleright ||||| | indicates concatenation operation
@@
We now investigate decidability problems for $\theta$-conjugate-freeness of $L$ when $L$ is either regular or context-free. Authors in [9] proved that it is decidable to determine whether or not a regular language contains two distinct conjugates of a word. We now prove that it is also decidable to determine whether or not a regular language $L$ with no $\theta$-palindrome contains two distinct $\theta$-conjugates of a word.
@@
Consider $L$ is a regular language with no $\theta$-palindromes. From Proposition 3.4, we know that $C_{\theta}(u^{\prime}a)=\{u^{\prime}a\}\cup\theta(a)C_{\theta}(u^{\prime})$ for $u^{\prime}\in\Sigma^{*}$ and $a\in\Sigma$. Then, $C_{\theta}(L)=\bigcup\limits_{w\in L}C_{\theta}(w)=L\cup\bigcup\limits_{\begin{subarray}{c}u^{\prime}a\in L\\ u^{\prime}\in\Sigma^{*},a\in\Sigma\end{subarray}}\theta(a)C_{\theta}(u^{\prime})=L\cup T$ where $T=\bigcup\limits_{\begin{subarray}{c}u^{\prime}a\in L\\ u^{\prime}\in\Sigma^{*},a\in\Sigma\end{subarray}}\theta(a)C_{\theta}(u^{\prime})$. Now, using a similar proof technique of Theorem 4.3, it is not hard to prove that if $L$ is regular, then $T$ is also regular. We now prove that $L$ is $\theta$-conjugate-free if and only if $L\cap T=\emptyset$. Consider $L$ is $\theta$-conjugate-free. Then for all $w\in L$, $C_{\theta}(w)\cap L=\{w\}$. If $x\in L\cap T$, then $x\in L$ and $x\in T$. This implies $x\in\theta(a)C_{\theta}(z)$ for some $za\in L$. If $x=za$, then for $\alpha\in\Sigma^{+}$ and $\beta\in\Sigma^{*}$, $x=\theta(\alpha)\beta$, $za=\beta\alpha$ and $\theta(\alpha)\beta=\beta\alpha$ which implies $x$ is a $\theta$-palindrome, a contradiction to the fact that $L$ does not contain any $\theta$-palindrome. Thus, $x\neq za$, i.e., $\{x,za\}\subseteq C_{\theta}(za)\cap L$ which implies $L$ is not $\theta$-conjugate-free, a contradiction. Hence, $L\cap T=\emptyset$. Now, consider $L\cap T=\emptyset$. If $w,w^{\prime}\in L$ such that $w\neq w^{\prime}$ and $w^{\prime}\in C_{\theta}(w)$, then $w^{\prime}\in T$ which implies $w^{\prime}\in L\cap T$, a contradiction. Thus, $L$ is $\theta$-conjugate-free. Therefore, $L$ is $\theta$-conjugate-free if and only if $L\cap T=\emptyset$. Now, $L$ and $T$ are regular languages. Since the emptiness of the intersection of two regular languages is decidable ([35]), we can decide whether or not L is $\theta$-conjugate-free. ∎
Now, we discuss the decidability result for a context-free language. First, we recall the following. An instance of the Post’s Correspondence Problem (PCP) [32] consists of a positive integer $n$, and two ordered $n$-tuples of strings $(U,V)$, where $U=(u_{0},u_{1},\ldots,u_{n-1})$ and $V=(v_{0},v_{1},\ldots,v_{n-1})$, $u_{i},v_{i}\in\Sigma^{*}$, $10\leq i\leq n-$ 1. A solution for the PCP instance $(U,V)$ is a sequence of integers $i_{1},i_{2},\ldots,i_{k}$, $10\leq i_{j}\leq n-$ 1, $j=1,2,\ldots,k$, $k\geq 1$, such that
@@
[32] The decision problem of determining whether or not a given PCP instance has a solution is undecidable.
@@
Let $\Sigma$ be alphabet and $(U,V)$ be an instance of PCP problem where $U=(u_{0},u_{1},\ldots,u_{n-1})$, $V=(v_{0},v_{1},\ldots,v_{n-1})$ with each $u_{i},v_{i}\in\Sigma^{*}$. Assume that the symbols #,\$$\#,\$,a_{0},a_{1},\cdots,a_{n-1}$ are distinct and not in $\Sigma$. Consider $\Sigma^{\prime}=\Sigma\cup\{\#,\$,a_{0},a_{1},\cdots,a_{n-1}\}$. Let $\theta:\Sigma^{\prime*}\rightarrow\Sigma^{\prime*}$ be an antimorphic involution defined as
  1. (i)
@@
($B$) \$$\$a_{j_{1}}a_{j_{2}}\cdots a_{j_{l}}\theta(v_{j_{l}})\cdots\theta(v_{j_{2}})\theta(v_{j_{1}})\$$
Under our assumed conditions on #,\$$\#,\$,a_{0},a_{1},\cdots,a_{n-1},\theta,u_{0},u_{1},\cdots u_{n-1},v_{0},v_{1},\cdots,v_{n-1}$, we observe that the symbol ##\## appears only at the left and right ends of words in the form $A$, the symbol \$currency-dollar\\$\$ does not appears in words of the form ${A}$, the symbol \$currency-dollar\\$\$ appears only at the left and right ends of words in the form ${B}$, and the symbol ##\## does not appears in words of the form ${B}$. We now show that $L(G)$ is not $\theta$-conjugate-free if and only if the PCP instance $(U,V)$ has a solution.
@@
$\beta^{\prime}=\$a_{i_{0}}a_{i_{1}}\cdots a_{i_{k}}\theta(v_{i_{k}})\cdots\theta(v_{i_{1}})\theta(v_{i_{0}})\$=\$a_{i_{0}}a_{i_{1}}\cdots a_{i_{k}}\theta(v_{i_{0}}v_{i_{1}}\cdots v_{i_{k}})\$=\$a_{i_{0}}a_{i_{1}}\cdots a_{i_{k}}\theta(u_{i_{0}}u_{i_{1}}\cdots u_{i_{k}})\$$
@@
[…]
@@
    As $\theta(\#)=\$$ and $y\in{\Sigma^{\prime}}^{+}$, $\theta(y)x$ starts with \$currency-dollar\\$\$. This implies $\theta(y)x$ must be in the form $B$. Then for some integer $r$,
@@
    \$$\$a_{j_{1}}a_{j_{2}}\cdots a_{j_{r}}\theta(v_{j_{r}})\cdots\theta(v_{j_{2}})\theta(v_{j_{1}})\$=\theta(\#u_{i_{1}}u_{i_{2}}\cdots u_{i_{m}}a_{i_{m}}\cdots a_{i_{2}}a_{i_{1}}\#)$
    $\implies$ \$$\$a_{j_{1}}a_{j_{2}}\cdots a_{j_{r}}\theta(v_{j_{r}})\cdots\theta(v_{j_{2}})\theta(v_{j_{1}})\$=\$a_{i_{1}}a_{i_{2}}\cdots a_{i_{m}}\theta(u_{i_{m}})\cdots\theta(u_{i_{2}})\theta(u_{i_{1}})\$.$
@@
Therefore, as PCP is undecidable [32], determining whether or not $L$ is $\theta$-conjugate-free is also undecidable when $L$ is context-free. ∎
7. Conclusions

### v5

@@ 5 Iterated $\theta$-conjugate of Words And Languages
Explanation of Step 2 through an Example: After Step 1, we have $w=C_{a_{1}}C_{b_{1}}C_{a_{1}}C_{\theta(a_{2})}$. Then, $S=\{C_{a_{1}}C_{b_{1}}C_{a_{1}}C_{\theta(a_{2})},C_{a_{1}}C_{b_{1}}C_{\theta(a_{2})}C_{a_{1}},C_{a_{1}}C_{a_{1}}C_{b_{1}}C_{\theta(a_{2})},C_{a_{1}}C_{a_{1}}C_{\theta(a_{2})}C_{b_{1}},C_{a_{1}}C_{\theta(a_{2})}C_{a_{1}}C_{b_{1}},\\ C_{a_{1}}C_{\theta(a_{2})}C_{b_{1}}C_{a_{1}},C_{b_{1}}C_{a_{1}}C_{a_{1}}C_{\theta(a_{2})},C_{b_{1}}C_{a_{1}}C_{\theta(a_{2})}C_{a_{1}},C_{b_{1}}C_{\theta(a_{2})}C_{a_{1}}C_{a_{1}},C_{\theta(a_{2})}C_{a_{1}}C_{a_{1}}C_{b_{1}},\\ C_{\theta(a_{2})}C_{a_{1}}C_{b_{1}}C_{a_{1}},C_{\theta(a_{2})}C_{b_{1}}C_{a_{1}}C_{a_{1}}\}$. Using multinomial coefficients, $|S|={4\choose 2~1~1}=12$.
@@
We now show that $R=C_{\theta}^{*}(w)$. Consider $u\in R$. Then by the above construction, $|u|_{a,\theta(a)}=|w|_{a,\theta(a)}$ for all $a\in\Sigma$. Thus, $u\in C_{\theta}^{*}(w)$. Therefore, $R\subseteq C_{\theta}^{*}(w)$. Consider $v\in C_{\theta}^{*}(w)$. Then, $|v|_{a,\theta(a)}=|w|_{a,\theta(a)}$ for all $a\in\Sigma$. Now, we replace each $a_{i}$ and $\theta(a_{i})$ in $v$ and $w$ by $C_{a_{i}}$ where $1\leq i\leq l$. Also, we replace each $b_{j}$ in $v$ and $w$ by $C_{b_{j}}$ where $1\leq j\leq r$. After replacements, consider $v$ and $w$ become $v^{\prime}$ and $w^{\prime}$, respectively, over $C_{1}\cup D$. Since $|v|_{a,\theta(a)}=|w|_{a,\theta(a)}$ for all $a\in\Sigma$, $v^{\prime}$ is a permutation of $w^{\prime}$, i.e., $v^{\prime}\in S$. Then by step 3, $v\in R$. Thus, $C_{\theta}^{*}(w)\subseteq R$. Therefore, $R=C_{\theta}^{*}(w)$. ∎
@@
  1. 1.
    Consider $w=cag$. Then $A_{1}=\{a,c,g\}$, $A_{2}=\emptyset$, $B=\emptyset$, $|w|_{a,b}=n_{a}=1$, $|w|_{c,d}=n_{c}=1$, $|w|_{g,h}=n_{g}=1$, and $n_{a}+n_{c}+n_{g}=n^{\prime}=3$. Thus by Theorem 5.9, $|{C_{\theta}}^{*}(w)|={3\choose 1\;1\;1}\times 2^{3}$ =48=48. In Example 5.8, we have explicitly computed the set ${C_{\theta}}^{*}(cag)$. From this set, we can verify that $|{C_{\theta}}^{*}(cag)|=48$.
  2. 2.
    Consider $w=caa$. Then, $A_{1}=\{a,c\}$, $A_{2}=\emptyset$, $B=\emptyset$, $|w|_{c,d}=n_{c}=1$, $|w|_{a,b}=n_{a}=2$, and $n_{c}+n_{a}=n^{\prime}=3$. Thus by Theorem 5.9, $|{C_{\theta}}^{*}(w)|={3\choose 1\;2}\times 2^{3}$ =24=24. In Example 5.8, we have explicitly computed the set ${C_{\theta}}^{*}(caa)$. From this set, we can verify that $|{C_{\theta}}^{*}(caa)|=24$.
  3. 3.
    Consider $w=aee$. Then, $A_{1}=\{a\}$, $A_{2}=\emptyset$, $B=\{e\}$, $|w|_{a,b}=n_{a}=1$, $|w|_{e}=m_{e}=2$, and $n_{a}=n^{\prime}=1$. Thus by Theorem 5.9, $|{C_{\theta}}^{*}(w)|={3\choose 1\;2}\times 2^{1}$ =6=6. We now verify the cardinality of ${C_{\theta}}^{*}(w)$ by computing it. Here, $C_{\theta}(w)=\{eeb,eea,eae,aee\}$, ${C_{\theta}}^{2}(w)=\{eeb,eea,bee,eae,ebe,aee\}$, and ${C_{\theta}}^{i}(w)={C_{\theta}}^{2}(w)$ for all $i\geq 2$. Thus, ${C_{\theta}}^{*}(w)={C_{\theta}}^{2}(w)$ and $|{C_{\theta}}^{*}(w)|=6$.
@@
It can be observed from Remark 5.11 that if $L$ is finite, then ${C_{\theta}}^{i}(L)$ is finite for $i\geq 0$. Also, if ${C_{\theta}}^{i}(L)$ is finite for some $i\geq 0$, then as $\subseteq{C_{\theta}}^{i}(L), LL$ is finite. This implies that $L$ is finite if and only if ${C_{\theta}}^{i}(L)$ is finite for $i\geq 0$.
@@
Consider a regular language $L=\{(bce)^{*}\}$ over $\{b,c,d,e,f\}$ where $\theta(b)=b$, $\theta(c)=d$ and $\theta(e)=f$. Since ${C_{\theta}}^{*}(w)=\{u\;|\;|u|_{a,\theta(a)}=|w|_{a,\theta(a)}~\forall a\in\Sigma\}$, ${C_{\theta}}^{*}(L)=\bigcup\limits_{w\in L}\{u\;|\;|u|_{a,\theta(a)}=|w|_{a,\theta(a)}~\forall a\in\Sigma\}.$ Then, ${C_{\theta}}^{*}(L)\cap\{b^{m_{1}}c^{m_{2}}e^{m_{3}}\;|\;m_{i}\geq 0\}=\{b^{n}c^{n}e^{n}\;|\;n\geq 0\}$, which is not context-free (using Theorem 4.4). Note that $\{b^{m_{1}}c^{m_{2}}e^{m_{3}}\;|\;m_{i}\geq 0\}$ is a regular language. Since the regular languages and the context-free languages are closed under intersection with regular languages (Shallit (2008)), the claim holds. ∎
6 $\theta$-conjugate-free Languages
@@
6.1 Decidability of $\theta$-conjugate-freeness
@@
Define an empty array AA of size $n$
@@
  for $l=n$ to 11 do $z\leftarrow z~||~\theta(u[l])$ $\triangleright$ $||$ indicates concatenation operation
  end for
end if
@@
We now investigate decidability problems for $\theta$-conjugate-freeness of $L$ when $L$ is either regular or context-free. Endrullis et al. (2017) proved that it is decidable to determine whether or not a regular language contains two distinct conjugates of a word. We now prove that it is also decidable to determine whether or not a regular language $L$ with no $\theta$-palindrome contains two distinct $\theta$-conjugates of a word.
@@
Consider $L$ is a regular language with no $\theta$-palindromes. From Proposition 3.4, we know that $C_{\theta}(u^{\prime}a)=\{u^{\prime}a\}\cup\theta(a)C_{\theta}(u^{\prime})$ for $u^{\prime}\in\Sigma^{*}$ and $a\in\Sigma$. Then, $C_{\theta}(L)=\bigcup\limits_{w\in L}C_{\theta}(w)=L\cup\bigcup\limits_{\begin{subarray}{c}u^{\prime}a\in L\\ u^{\prime}\in\Sigma^{*},a\in\Sigma\end{subarray}}\theta(a)C_{\theta}(u^{\prime})=L\cup T$ where $T=\bigcup\limits_{\begin{subarray}{c}u^{\prime}a\in L\\ u^{\prime}\in\Sigma^{*},a\in\Sigma\end{subarray}}\theta(a)C_{\theta}(u^{\prime})$. Now, using a similar proof technique of Theorem 4.3, it is not hard to prove that if $L$ is regular, then $T$ is also regular. We now prove that $L$ is $\theta$-conjugate-free if and only if $L\cap T=\emptyset$. Consider $L$ is $\theta$-conjugate-free. Then for all $w\in L$, $C_{\theta}(w)\cap L=\{w\}$. If $x\in L\cap T$, then $x\in L$ and $x\in T$. This implies $x\in\theta(a)C_{\theta}(z)$ for some $za\in L$. If $x=za$, then for $\alpha\in\Sigma^{+}$ and $\beta\in\Sigma^{*}$, $x=\theta(\alpha)\beta$, $za=\beta\alpha$ and $\theta(\alpha)\beta=\beta\alpha$ which implies $x$ is a $\theta$-palindrome, a contradiction to the fact that $L$ does not contain any $\theta$-palindrome. Thus, $x\neq za$, i.e., $\{x,za\}\subseteq C_{\theta}(za)\cap L$ which implies $L$ is not $\theta$-conjugate-free, a contradiction. Hence, $L\cap T=\emptyset$. Now, consider $L\cap T=\emptyset$. If $w,w^{\prime}\in L$ such that $w\neq w^{\prime}$ and $w^{\prime}\in C_{\theta}(w)$, then $w^{\prime}\in T$ which implies $w^{\prime}\in L\cap T$, a contradiction. Thus, $L$ is $\theta$-conjugate-free. Therefore, $L$ is $\theta$-conjugate-free if and only if $L\cap T=\emptyset$. Now, $L$ and $T$ are regular languages. Since the emptiness of the intersection of two regular languages is decidable (Shallit (2008)), we can decide whether or not L is $\theta$-conjugate-free. ∎
Now, we discuss the decidability result for a context-free language. First, we recall the following. An instance of the Post’s Correspondence Problem (PCP) Post (1946) consists of a positive integer $n$, and two ordered $n$-tuples of strings $(U,V)$, where $U=(u_{0},u_{1},\ldots,u_{n-1})$ and $V=(v_{0},v_{1},\ldots,v_{n-1})$, $u_{i},v_{i}\in\Sigma^{*}$, $0\leq i\leq n-1$. A solution for the PCP instance $(U,V)$ is a sequence of integers $i_{1},i_{2},\ldots,i_{k}$, $0\leq i_{j}\leq n-1$, $j=1,2,\ldots,k$, $k\geq 1$, such that
@@
Post (1946) The decision problem of determining whether or not a given PCP instance has a solution is undecidable.
@@
Let $\Sigma$ be alphabet and $(U,V)$ be an instance of PCP problem where $U=(u_{0},u_{1},\ldots,u_{n-1})$, $V=(v_{0},v_{1},\ldots,v_{n-1})$ with each $u_{i},v_{i}\in\Sigma^{*}$. Assume that the symbols $a_{0},a_{1},\cdots,a_{n-1}$ are distinct and not in $\Sigma$. Consider $\Sigma^{\prime}=\Sigma\cup\{\#,\$,a_{0},a_{1},\cdots,a_{n-1}\}$. Let $\theta:\Sigma^{\prime*}\rightarrow\Sigma^{\prime*}$ be an antimorphic involution defined as
  1. 1.
@@
$a_{j_{1}}a_{j_{2}}\cdots a_{j_{l}}\theta(v_{j_{l}})\cdots\theta(v_{j_{2}})\theta(v_{j_{1}})\$$ ($B$)
Under our assumed conditions on $a_{0},a_{1},\cdots,a_{n-1},\theta,u_{0},u_{1},\cdots u_{n-1},v_{0},v_{1},\cdots,v_{n-1}$, we observe that the symbol #\# appears only at the left and right ends of words in the form $A$, the symbol \$\\$ does not appears in words of the form A{A}, the symbol \$\\$ appears only at the left and right ends of words in the form B{B}, and the symbol #\# does not appears in words of the form B{B}. We now show that $L(G)$ is not $\theta$-conjugate-free if and only if the PCP instance $(U,V)$ has a solution.
@@
$\beta^{\prime}=\$a_{i_{0}}a_{i_{1}}\cdots a_{i_{k}}\theta(v_{i_{k}})\cdots\theta(v_{i_{1}})\theta(v_{i_{0}})\$$ $=\$a_{i_{0}}a_{i_{1}}\cdots a_{i_{k}}\theta(v_{i_{0}}v_{i_{1}}\cdots v_{i_{k}})\$$
$=\$a_{i_{0}}a_{i_{1}}\cdots a_{i_{k}}\theta(u_{i_{0}}u_{i_{1}}\cdots u_{i_{k}})\$$
@@
As $\theta(\$)=\#$ and $\theta(c)=c~\forall c\in\{a_{0},a_{1},\cdots,a_{n-1}\}$, we have $\beta^{\prime}=\theta(\alpha^{\prime})$. Since $\neq\$$, $\alpha^{\prime}\neq\beta^{\prime}$. Thus, $L(G)$ is not $\theta$-conjugate-free.
@@
    As $\theta(\#)=\$$ and $y\in{\Sigma^{\prime}}^{+}$, $\theta(y)x$ starts with \$\\$. This implies $\theta(y)x$ must be in the form $B$. Then for some integer $r$,
@@
    $\$a_{j_{1}}a_{j_{2}}\cdots a_{j_{r}}\theta(v_{j_{r}})\cdots\theta(v_{j_{2}})\theta(v_{j_{1}})\$=\theta(\#u_{i_{1}}u_{i_{2}}\cdots u_{i_{m}}a_{i_{m}}\cdots a_{i_{2}}a_{i_{1}}\#)$
    $\implies$ $\$a_{j_{1}}a_{j_{2}}\cdots a_{j_{r}}\theta(v_{j_{r}})\cdots\theta(v_{j_{2}})\theta(v_{j_{1}})\$=\$a_{i_{1}}a_{i_{2}}\cdots a_{i_{m}}\theta(u_{i_{m}})\cdots\theta(u_{i_{2}})\theta(u_{i_{1}})\$.$
@@
Therefore, as PCP is undecidable Post (1946), determining whether or not $L$ is $\theta$-conjugate-free is also undecidable when $L$ is context-free. ∎
7 Conclusions